12 ml of a 51% acid solution was mixed with 4ml of a 87% acid solution. Find the concentration of the new mixture.
Found 2 solutions by mananth, stanbon:
Answer by mananth(16946)

12 ml of a 51% acid solution was mixed with 4ml of a 87% acid solution. Find the concentration of the new mixture.
12*51%+4*87%=16*x%
multiply by 100
12*51+4*87=16x
16x=960
/16
x=60%

Answer by stanbon(75887)

12 ml of a 51% acid solution was mixed with 4ml of a 87% acid solution. Find the concentration of the new mixture.
------------------
Equation:
acid + acid = acid
0.51*12 + 0.87*4 = x(12+4)
----
51*12 + 87*4 = x(1600)
----
x = 0.60 = 60%
-------------------
Cheers,
Stan H.
---------------------